what is the solution in this question the length of the rectangle is 2 times it's width.the area of the rectangle is 72 squares inches. find the dimension of the rectangle
Answer by graphmatics(170)

Let w be the width of the rectangle then the length of the rectangle is 2w. The area of the rectangle is the width times the length so w*2w=72 so w^2=72/2 so w^2=36 so w=sqrt(36). w=width=6. l=length=12.